What is the frequency of a photon with a wavelength of 781 nm. Report your answers to three significant digits.

The frequency is = _________ s-1

1
Expert's answer
2021-07-12T02:59:43-0400

The frequency ƒ = the speed of light c / the wavelenght λ

ƒ = 3.0×108 m/s × 7.81×10-7 m = 3.84 ×1014 Hz

Answer: "\u0192 = 3.84 \u00d710^{14} Hz"